OpenStudy (bananas1234):

2. Simplify. a) (3xy3)(2x3y) b) 2y(2y + 3y2) – y2(3 – y)

OpenStudy (luigi0210):

Try distributing

OpenStudy (michele_laino):

question b) we can write this \[\begin{gathered} 2y\left( {2y + 3{y^2}} \right) - {y^2}\left( {3 - y} \right) = \hfill \\ = \left( {2y \cdot 2y} \right) + \left( {2y \cdot 3{y^2}} \right) + \left( { - {y^2}} \right) \cdot 3 + \left( { - {y^2}} \right) \cdot \left( { - y} \right) = ...? \hfill \\ \end{gathered} \]

jhonyy9 (jhonyy9):

do you understand it now ?

jhonyy9 (jhonyy9):

a) multiply the same terms 3*2 =6 and x*x^3 = x^(1+3) = x^4 and y^3 *y = y^(3+1) = y^4 hope this is understandably more easy so try the second like first ok ?

jhonyy9 (jhonyy9):

hey !!! are you here ?

jhonyy9 (jhonyy9):

so then you dont like cooperate nothing bye !

OpenStudy (bananas1234):

sorry i had to go. @jhonyy9